# Variability of the UV luminosity function with SPICE

### v1

@@ Acknowledgements
All the analysis were carried out on the machines of Max Planck Institute for Astrophysics (MPA) and Max Planck Computing and Data Facility (MPCDF). AB thanks the entire EoR research group of MPA for all the encouraging comments for this project. This work made extensive use of publicly available software packages : numpy (van der Walt et al., 2011), matplotlib (Hunter, 2007), scipy (Jones et al., 2001). Authors thank the developers of these packages.
@@
Appendix A Comparison between different methods for computing $\sigma\rm{{}_{UV}}$
@@
In this section, we discuss how our major findings are affected by adopting the method proposed by Pallottini & Ferrara (2023) (hereafter referred to as PF23) in the calculation of the variability. We begin by analyzing the temporal evolution of the SFR for the most massive halo at $z=5$ illustrated in Figure 4. Each curve is fitted with a third-order polynomial on a log-linear scale (i.e. $\text{SFR}_{\text{fit}}$), following the technique used by PF23. The stochastic variability of the SFR is then quantified by computing the residuals, defined as $\delta_{\rm{SFR}}=\log_{10}(\text{SFR}/\text{SFR}_{\text{fit}})$, from which we derive the standard deviation, $\rm{\mathit{\sigma}_{SFR}}$. The results of this approach are compared with the standard method presented in Section 3.2 and summarized in Figure 11. We observe that the $\delta_{\rm{SFR}}$ distribution using the PF23 method is slightly shifted toward higher values across all models compared to the standard method. This shift is reflected in the slightly higher values of $\rm{\mathit{\sigma}_{SFR}}$, with 0.39 for bursty-sn, 0.17 for smooth-sn, and 0.19 for hyper-sn, each exceeding the standard method’s results by approximately $\Delta(\mathit{\sigma}_{\rm{SFR}})\sim 0.04-0.05$. To examine the behavior of the entire halo population at $z=5$, Figure 12 presents the distribution of $\sigma_{\mathrm{SFR}}$ calculated using both methods across all SN feedback models. Interestingly, we observe that the PF23 method yields a distribution of $\sigma_{\mathrm{SFR}}$ values that are slightly lower across all models when compared to the standard method. This finding contrasts with the trend seen in Figure 11. Specifically, we note that the peak of the $\sigma_{\mathrm{SFR}}$ distribution shifts by approximately $0.15-0.2$.
@@
Appendix B Fitting and Extrapolation of UVLF
To investigate whether the SPICE UVLF, incorporating scatter, can effectively reproduce the observed luminosity functions, we fit each UVLF with a Schechter (1976) function [^2] The functional form is given by :
$\rm{\Phi(\mathit{M}_{UV})=0.4\ln(10)\,\phi^{*}\,\left[10^{0.4(\mathit{M}_{UV}^{*}-\mathit{M}_{UV})(\alpha+1)}\right]\exp\left[-10^{0.4(\mathit{M}_{UV}^{*}-\mathit{M}_{UV})}\right]}$, where $\phi^{*}$ is the normalization factor, $\rm{\mathit{M}_{UV}^{*}}$ is the characteristic magnitude and $\alpha$ is the faint-end slope.
. In Table 1, we show the fitted parameters for $z=5$, 6, 7, 9, 10, 12. We note that for $\rm{\mathit{M}_{UV}\lesssim-19}$, SPICE lacks a sufficient statistical sample of halos for a robust fitting. As a result, the Schechter function fails to provide an optimal fit, leading to progressively higher deviations from the SPICE LFs towards lower $\rm{\mathit{M}_{UV}}$ values relative to the fitted characteristic magnitude ($\rm{\mathit{M}_{UV}^{*}}$). Consequently, the reduced $\rm{\chi^{2}}$ values are notably low in some cases where the deviations are larger, particularly at higher redshifts.
As a reference, in Figure 13 we plot the fitted curves at $z=10$ and 12 as solid lines, while the shaded regions represent the 2-$\sigma$ scatter obtained by performing the fits for the curves corresponding to 10, 30, and 50 Myr before and after the reference redshift. As mentioned above, the fit at the brightest end of the LF is not robust, and indeed all models fall below observational data, although it has previously been noticed (Bowler et al., 2020; Donnan et al., 2024; Whitler et al., 2025) that high-$z$ galaxies with $\rm{\mathit{M}_{UV}\lesssim-22}$ are in excess of the exponential decline of the Schechter function. At both redshifts, the bursty-sn and smooth-sn models perform similarly in reproducing observations within the range $-19\lesssim\rm{\mathit{M}_{UV}}\lesssim-17$, with the latter exhibiting slightly smaller scatter. Both models align more closely with the data toward the brighter end, up to $\rm{\mathit{M}_{UV}}\approx-20$, at $z\approx 12$. Beyond this point, they systematically underpredict the data, with deviations becoming progressively larger at lower $\rm{\mathit{M}_{UV}}$. The hyper-sn model, while displaying a greater scatter in comparison to smooth-sn at $z\approx 12$, does not achieve better agreement with the data, instead showing a significant offset from observations. At $z\approx 10$, however, its predictions remain comparable to those of other models, with the lowest scatter observed. To provide additional information, we also include the 2-$\sigma$ scatter obtained by linearly extrapolating the SPICE UVLFs corresponding to 10, 30, and 50 Myr before and after the reference redshift beyond the $\rm{\mathit{M}_{UV}}$ range available in SPICE, without applying any fitting procedure. These curves align more closely with observations across the entire $\mathit{M}_{\rm UV}$ range, with hyper-sn offering the best overall agreement at both redshifts. The bursty-sn model tends to overpredict the counts at $z\approx 12$, but closely matches most observational constraints at $z\approx 10$. Conversely, smooth-sn performs the worst overall, but shows better agreement at $z\approx 12$ than at $z\approx 10$.
Figure 13: UVLF at $z=12$ (top panel) and 10 (bottom) obtained after fitting with a Schechter (1976) function (solid lines). Shaded regions represent the 2-$\sigma$ scatter calculated from the evolution of the UVLF over time intervals of 10, 30, and 50 Myr around the reference redshifts. The hatched shaded regions indicate a direct extrapolations of the UVLFs. Colors refer to different SN feedback models. The black dashed curve indicates the result from Whitler et al. (2025) fitted to their observational F115W dropout sample assuming a Schechter (1976) function, while a compilation of observations from HST and JWST (Bouwens et al., 2015; Harikane et al., 2022; Naidu et al., 2022; Adams et al., 2023; Harikane et al., 2023; Bouwens et al., 2023a; Bouwens et al., 2023b; Leung et al., 2023; Donnan et al., 2023b; Donnan et al., 2023c; Pérez-González et al., 2023; Casey et al., 2024; Robertson et al., 2024; McLeod et al., 2024; Whitler et al., 2025) is shown as gray data points.
$z$ $\phi^{*}(\times 10^{-3})$ $\rm{\mathit{M}_{UV}^{*}}$ $\alpha$ $\rm{\chi_{\nu}^{2}}$
Bursty Smooth Hyper Bursty Smooth Hyper Bursty Smooth Hyper Bursty Smooth Hyper
5 $75.95$ $256.11$ $23.36$ $-17.32$ $-16.84$ $-17.81$ $-1.53$1.40 -1.67 $0.40$ 2.68 0.39
7 $44.19$ $155.57$ $63.89$ $-17.12$ $-16.62$ $-16.93$ $-1.61$1.45 -1.59 $0.45$ 1.92 0.78
9 $0.18$ $136.72$ $34.28$ $-19.96$ $-15.93$ $-16.53$ $-1.89$1.39 -1.65 $0.06$ 0.33 0.12
10 $30.84$ $67.20$ $12.23$ $-16.51$ $-16.14$ $-16.51$ $-1.48$1.48 -1.67 $0.21$ 0.15 0.06
12 $14.18$ $21.13$ $4.06$ $-16.38$ $-16.66$ $-15.22$ $-1.43$1.42 -1.55 $0.02$ 0.02 0.01
Table 1: Fitted parameters of the UVLF functional forms at various redshifts, modeled using the Schechter (1976) function. The parameters include the overall normalization ($\phi^{*}$), the characteristic magnitude ($\rm{\mathit{M}_{UV}^{*}}$), and the faint-end slope ($\alpha$). The reduced chi-square ($\chi^{2}_{\nu}$) values assess the goodness of fit.

### v2

@@ Acknowledgements
All the analysis were carried out on the machines of Max Planck Institute for Astrophysics (MPA) and Max Planck Computing and Data Facility (MPCDF). We thank the anonymous reviewer and the editor for the useful comments which helped to improve the manuscript. AB thanks the entire EoR research group of MPA for all the encouraging comments for this project. This work made extensive use of publicly available software packages : numpy (van der Walt et al., 2011), matplotlib (Hunter, 2007), scipy (Jones et al., 2001). Authors thank the developers of these packages.
@@
Appendix A Comparison between different methods for computing $\sigma\rm{{}_{SFR}}$
@@
In this section, we discuss how our major findings are affected by adopting the method proposed by Pallottini & Ferrara (2023) (hereafter referred to as PF23) in the calculation of the variability. We begin by analyzing the temporal evolution of the SFR for the most massive halo at $z=5$ illustrated in Figure 4. Each curve is fitted with a third-order polynomial on a log-linear scale (i.e. $\text{SFR}_{\text{fit}}$), following the technique used by PF23. The stochastic variability of the SFR is then quantified by computing the residuals, defined as $\delta_{\rm{SFR}}=\log_{10}(\text{SFR}/\text{SFR}_{\text{fit}})$, from which we derive the standard deviation, $\rm{\mathit{\sigma}_{SFR}}$. The results of this approach are compared with the standard method presented in Section 3.2 and summarized in Figure 11. We observe that the $\delta_{\rm{SFR}}$ distribution using the PF23 method is slightly shifted toward higher values across all models compared to the standard method. This shift is reflected in the slightly higher values of $\rm{\mathit{\sigma}_{SFR}}$, with 0.39 for bursty-sn, 0.17 for smooth-sn, and 0.19 for hyper-sn, each exceeding the standard method’s results by approximately $\Delta(\mathit{\sigma}_{\rm{SFR}})\sim 0.04-0.05$. To examine the behavior of the entire halo population at $z=5$, Figure 12 presents the distribution of $\sigma_{\mathrm{SFR}}$ calculated using both methods across all SN feedback models. Interestingly, we observe that the PF23 method yields a distribution of $\sigma_{\mathrm{SFR}}$ values that are slightly lower across all models when compared to the standard method. This finding contrasts with the trend seen in Figure 11. Specifically, we note that the peak of the $\sigma_{\mathrm{SFR}}$ distribution shifts by approximately 0.150.20.15-0.2.